OpenStudy (babynini):

Use traces to identify the shape x^2-y^2+z^2=1

OpenStudy (babynini):

So I know about plugging k or 0 in to try and identify each plane, but I get weird equations on this one and can't identify the shape

OpenStudy (babynini):

@zepdrix :)

OpenStudy (babynini):

yz plane, x=k is x^2-y^2=1-k^2

OpenStudy (babynini):

or putting x=0, z^2-y^2=1

OpenStudy (babynini):

which would be a hyperbola, correct?

OpenStudy (babynini):

similarily z=0, gives us x^2-y^2=1 which is another hyperbola y=0, x^2+z^2=1 is a circle

OpenStudy (babynini):



OpenStudy (babynini):

hahah well something like this, only less potato-y!